# Louis Pallegoix1, Jaime Travesedo1, Alexandre S. May1,2, Léo Balembois1, Denis Vion1, Patrice Bertet*1 and Emmanuel Flurin Université Paris-Saclay, CEA, CNRS, SPEC, 91191 Gif-sur-Yvette Cedex, France 2Alice&\&&Bob, 53 boulevard du Général Martial Valin, 75015 Paris

### v1

@@ Abstract
We report on the characteristics of a microwave photon counter device based on a superconducting transmon qubit. Its design is similar to [1], with an additional bandwidth tuning circuit that allows optimizing the device efficiency and noise. Owing to this new feature and to improvements in device fabrication, a power sensitivity of $3\cdot 10^{-23}\ \mathrm{W}/\sqrt{\mathrm{Hz}}$ is reached. We confirm the high performance of the device by measuring single spin microwave fluorescence.
@@ I Introduction
Single-photon detection has long been a fundamental tool in quantum technologies, with well-established applications in the optical domain, ranging from fluorescence microscopy [2, 3, 4, 5] to quantum communication [6]. However, extending single-photon detection into the microwave regime presents significant challenges. The energy of microwave photons—on the order of µeV—is several orders of magnitude lower than that of optical photons, requiring detectors to operate at cryogenic temperatures to suppress thermal noise and reduce the background thermal photon population.
Despite these challenges, Single Microwave Photon Detectors (SMPDs) have emerged as critical tools for advancing quantum sensing and information processing. Their ability to detect individual microwave photons has opened new possibilities for a wide range of applications, including the detection of weak incoherent microwave emitters, such as electron spins in solids [7, 8, 9], and even searches for hypothetical dark matter particles [10, 11]. SMPDs also hold potential for primary thermometry at low temperatures [12] and for use in quantum illumination protocols [13]. Moreover, these detectors can play a key role in quantum information processing [14, 15, 16], enabling heralded entanglement between superconducting qubits [17], improving qubit readout schemes [18], and facilitating the generation of robust quantum states [19].
Various SMPD designs have been proposed and implemented, with most relying on superconducting qubits [20, 7, 17, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32], quantum dots [33, 34] or bolometric detectors [35]. These devices have demonstrated promising results, achieving sensitivities in the range of $10^{-21}$ to $10^{-22}\,\mathrm{W/\sqrt{Hz}}$ at microwave frequencies. However, there remain limitations in terms of bandwidth, detection efficiency, and dark count rates that must be addressed to further enhance their performance.
In this work, we present an enhanced version of the SMPD introduced in Ref. [1], based on a superconducting qubit and a four-wave mixing process. The new design integrates a frequency-tunable Purcell filter, providing control over the detection bandwidth. This tunability significantly reduces thermal noise and lowers the dark count rate down to $30\ \mathrm{s}^{-1}$, improving the overall detector performance. Additionally, advancements in the fabrication process have led to an extended qubit relaxation time ($T_{1}$), which in turn enhances the detection efficiency to $0.8$. As a result of these improvements, the detector achieves a power sensitivity of $3\cdot 10^{-23}\,\mathrm{W/\sqrt{Hz}}$, representing a significant advancement over previous designs. We confirm the SMPD improved sensitivity by detecting the microwave fluorescence from a single spin.
@@ II.1 Device design
The SMPD device (see Fig. 1a and b) consists of a transmon qubit (mode $q$, frequency $\omega_{q}$), capacitively coupled to two harmonic oscillator modes, called respectively the buffer (mode $b$, frequency $\omega_{b}$), and the waste (mode $w$, frequency $\omega_{w}$). The buffer is connected to the input line of the device through another resonator at frequency $\omega_{pb}$. The latter serves as Purcell filter to decouple the qubit from the input line, while also setting the effective buffer energy coupling rate $\kappa_{\mathrm{b,c}}$. The total energy loss rate of the buffer resonator is $\kappa_{b}=\kappa_{\mathrm{b,c}}+\kappa_{\mathrm{b,i}}$, with $\kappa_{\mathrm{b,i}}$ the internal losses of the buffer. The waste is connected to an output line through another Purcell filter, at frequency $\omega_{pw}$, with a rate $\kappa_{w}$. Other relevant parameters are the qubit energy relaxation time $T_{1}$ and equilibrium excited state population, $p_{\mathrm{th},q}$, which varies with the qubit temperature $T$ as $p_{\mathrm{th},q}=\frac{1}{e^{\frac{\hbar\omega_{q}}{k_{B}T}}-1}$. In this formula, $T$ is not necessarily the cryostat base temperature, as various heating effects (due to infra-red radiation in particular) may lead to a higher effective qubit temperature.
For optimal operation as an energy detector, both the frequency and bandwidth of the device need to be matched to the source of interest, be it a spin [1, 8] or a dark matter candidate [36]. Frequency tuning is achieved by inserting a SQUID in the buffer mode, resulting in a buffer frequency $\omega_{b}(\phi_{b})$ that depends on the flux threading the buffer SQUID loop, $\phi_{b}$. An asymmetric SQUID is purposely chosen to minimize sensitivity to flux noise. Bandwidth tuning is achieved by inserting a second SQUID in the Purcell filter mode, resulting in a buffer coupling rate $\kappa_{\mathrm{b,c}}(\phi_{\mathrm{pb}})$ that depends on the flux threading the Purcell SQUID loop, $\phi_{\mathrm{pb}}$. This SQUID is symmetric to maximize the bandwidth tuning range.
@@ II.2 Four-Wave Mixing
The working principle of photon detection relies on a four-wave mixing (4WM) process that maps the presence or absence of a photon on the state of the transmon qubit. The non-linear element is provided by the qubit Josephson junction. The 4WM is activated by a microwave signal (called the pump, frequency $\omega_{p}$) of amplitude $\xi_{0}$, applied to the qubit via a pump line connected capacitively to the qubit (see Fig. 1). When the pump frequency approximately satisfies the relation $\omega_{b}+\omega_{p}=\omega_{q}+\omega_{w}-\chi_{w}$, the dominant process is described by the Hamiltonian
@@
where $\chi_{b}$ (resp. $\chi_{w}$) is the qubit dispersive coupling to the buffer mode (resp. the waste mode) [20]. This Hamiltonian reversibly converts a buffer and pump photon into a qubit excitation and waste photon. Dissipation of the waste photon into the output line ensures that the process is irreversible; on the other hand, the qubit remains excited long enough to be readout subsequently, and can thus be used as a flag that frequency conversion occurred.
An approximate description is obtained by eliminating the qubit degree of freedom, yielding an effective Hamiltonian $H_{\mathrm{eff}}=g_{\mathrm{4WM}}bw^{\dagger}+h.c.$ with $g_{\mathrm{4WM}}=\sqrt{\chi_{b}\chi_{w}}\xi_{0}$. It describes the parametric coupling between the buffer and waste modes, similar to the Josephson Parametric Converter for instance. Assuming that $\kappa_{\mathrm{b,i}}\ll\kappa_{\mathrm{b,c}}$, the efficiency for frequency conversion under this process is $\eta_{\mathrm{4WM}}=4C/(1+C)^{2}$, where $C=\frac{4\chi_{b}\chi_{w}|\xi_{0}|^{2}}{\kappa_{w}\kappa_{b}}$ is the system cooperativity. Unit efficiency is reached when $C=1$. Physically, this occurs when the frequency conversion rate $4|g_{\mathrm{4WM}}|^{2}/\kappa_{w}$ is equal to the input resonator coupling rate, $\kappa_{b}$. This condition can always be achieved by adjusting the pump frequency and amplitude (see Appendix A for more details).
Figure 1: Overview of the device. a) Global layout of the SMPD. b) Lumped-element scheme of the SMPD. The colors of the different resonators are in accordance with their colors on the chip scheme. Note : The buffer SQUID is asymmetrical. c) Operation cycle of the SMPD. During the Detection step (duration $T_{d}$), the pump is switched on; a photon arriving in the buffer mode is converted into a photon in the waste which quickly decays, and a qubit excitation which decays in a characteristic time $T_{1}\gg T_{d}$. Dispersive readout is performed through the waste resonator in a time $T_{\mathrm{RO}}$. If the qubit is found in |$\ket{e}$, a $\pi$ pulse is applied, followed by qubit readout, and re-application of the reset if it is found in |$\ket{e}$. This conditional reset step lasts $T_{\mathrm{reset}}$. The whole cycle lasts $16.6$ µs (assuming one reset cycle).
@@ II.3 Cyclic operation
The SMPD is operated cyclically. A cycle consists of three successive steps: Detection, Readout, and Reset (see Fig. 1c. During the detection step, which lasts $T_{d}$, the pump drive is simply switched on, with suitable amplitude and frequency for maximum frequency conversion. In the next step, dispersive qubit readout is achieved via the waste resonator by measuring the reflected field quadratures of a microwave tone in a time $T_{\mathrm{RO}}$. We define the readout fidelity $\mathcal{F}_{\mathrm{RO}}<1$ as the probability to measure the qubit in the excited state after a $\pi$ pulse.
In the third step, the qubit is reset in its ground state by applying a $\pi$ pulse conditioned to the outcome of the previous readout being the excited state. This resets the qubit in its ground state efficiently, with a probability that can be even lower than the thermal equilibrium qubit excited state probability. In the following we will assume this reset step to be ideal. Note that the duration of the reset step $T_{\mathrm{reset}}$ varies from cycle to cycle due to its conditional nature. The detector duty cycle $\eta_{\mathrm{cycle}}$ is defined as the average value of $T_{d}/(T_{d}+T_{\mathrm{RO}}+T_{\mathrm{reset}})$.
@@ III Figures of merit
In this section, we describe in somewhat detail the processes that limit the device performance as an energy detector. This performance is captured by the power sensitivity $S$, defined as the power that can be detected with unit signal-to-noise ratio in $1$ s integration time:
@@
In this formula, $\eta_{\mathrm{SMPD}}$ is the SMPD overall detection efficiency, and $\alpha$ is the average number of counts per second when no signal is sent to the device. The latter can be separated in two contributions $\alpha=\alpha_{\mathrm{err}}+\alpha_{\mathrm{th}}$, $\alpha_{\mathrm{err}}$ being the false positive events, and $\alpha_{\mathrm{th}}$ the counts due the detection of thermal photons in the input line. In the following, we address each of these three quantities, with the goal of maximizing $\eta_{\mathrm{SMPD}}$, and minimizing $\alpha_{\mathrm{err}}$ and $\alpha_{\mathrm{th}}$.
@@ III.1 Detection efficiency
$\eta_{\mathrm{SMPD}}(\omega)=\eta_{\omega}\eta_{\mathrm{4WM}}\eta_{q}\mathcal{F}_{\mathrm{RO}}\eta_{\mathrm{cycle}},$ (3)
where the detector frequency response function is approximated by a Lorentzian of width $\kappa_{d}$ (the detector linedwidth) such that $\eta_{\omega}=\frac{1}{1+\left[\frac{2(\omega-\omega_{b})}{\kappa_{d}}\right]^{2}}$ The qubit efficiency $\eta_{q}=\frac{T_{1}}{T_{d}}(1-\mathrm{e}^{-\frac{T_{d}}{T_{1}}})$ can be approximated by $1-\frac{T_{d}}{2T_{1}}$ in the limit $T_{d}\ll T_{1}$ [1]. To maximize $\eta_{\mathrm{SMPD}}$, $T_{d}$ should be optimized in order to maximize the duty cycle while keeping the cycle time well below $T_{1}$ to avoid relaxation of the qubit before readout.
@@ III.2 Dark count rate
The dark count rate $\alpha_{\mathrm{err}}$ is the number of counts detected per unit of time although no photon was present in the input line. It has two contributions, $\alpha_{\mathrm{err}}=\alpha_{q}+\alpha_{p}$. First, the qubit can be found in the excited state despite the active reset procedure. This is due to the relaxation towards the thermal equilibrium value $p_{q,th}$, which takes place during the detection window with a characteristic time $T_{1}$, giving rise to $\alpha_{q}=p_{q,th}(T_{d}/T_{1})(1/(T_{d}+T_{\mathrm{RO}}+T_{\mathrm{reset}}))$ (assuming that the active reset works perfectly). The other contribution $\alpha_{p}$ is due to heating brought by the pump. Note that the readout errors causing the ground state to be erroneously detected as the excited state are negligible because we purposely choose a detection threshold that minimizes this effect.
@@ III.3 Thermal counts
Thermalizing the microwave field at millikelvin temperatures is well-known to be a difficult task, because of vanishing thermal conductivity and of the large attenuation needed in the input lines. This results in an effective field temperature often significantly larger than the cryostat base temperature, and in our case in the presence of a finite flux of thermal photons at the SMPD input causing a thermal count rate of thermal origin
[…]
where $\bar{n}_{\mathrm{th,b}}=\frac{1}{e^{\frac{\hbar\omega_{b}}{k_{B}T}}-1}$ the photon number per mode assuming that the field is at temperature $T$. Because $\kappa_{d}\ll\omega_{b}$, we get
@@
Owing to the tunable bandwidth of our device, we are here able to test the predicted proportionality between noise and detector bandwidth, as will be seen below.
@@ IV.1 Device characterization
The SMPD resonators are measured with a Vector Network Analyser (VNA). Each resonance is fitted in order to determine the resonator frequency and linewidth. The result of these fits are reported in tab. 1.
@@
Table 1: Measured resonator frequencies and linewidths. The Purcell filter of the buffer and the buffer itself are tunable over the indicated interval. The buffer linewidth depends on the the Purcell frequency, as shown on fig. 2.
@@
Figure 2: Tunability curves. a) Full blue circules are the measured frequency of the Purcell filter, $\omega_{pb}$, as a function of the flux bias sent to its SQUID $\Phi_{pb}$. Solid line is a fit assuming a symmetric SQUID. b) Solid circles are the measured frequency of the buffer resonator, $\omega_{b}$, as a function of the flux bias applied to its SQUID, $\Phi_{b}$, for several values of $\Phi_{pb}$ (blue is $-0.45\Phi_{0}$, orange is $-0.24\Phi_{0}$, green is $-0.03\Phi_{0}$). The solid lines are sinusoidal fits, yielding the asymmetry between the buffer SQUID junctions. c) Full circles show the measured coupling rate of the buffer resonator to the $50\Omega$ line, as a function of its detuning from the Purcell filter, $\omega_{b}-\omega_{pb}$, for $3$ different values of $\Phi_{b}$ : blue for $\Phi_{b}=0$, orange for $\Phi_{b}=0.25\Phi_{0}$, green for $\Phi_{b}=0.5\Phi_{0}$. The dashed line shows the buffer internal losses, $\kappa_{b,i}$. d) Full circles are measured SMPD efficiency, $\eta_{\mathrm{SMPD}}$, as a function of signal frequency $\omega$, for $\kappa_{b}/2\pi=170$ kHz (blue) and $280$ kHz (orange). Dashed lines are Lorentzian fits yielding the SMPD bandwidth $\kappa_{d}/2\pi$, equal to respectively $314$ kHz and $531$ kHz.
@@
Figure  2a, shows $\omega_{b}(\phi_{b})$, for several values of $\phi_{pb}$. The buffer tuning range is narrow (60 MHz) and the curve is well fitted by a sinusoid [37], from which we deduce a ratio of 15 between the Josephson junction critical currents of the asymmetric SQUID. Little cross-talk between the $\phi_{b}$ and $\phi_{pb}$ is observed, as evidenced by the weak dependence of $\omega_{b}/2\pi$ on $\phi_{pb}$.
@@
From the fits of the buffer resonance curves, we extract $\kappa_{b,i}=2.2\cdot 10^{5}\mathrm{s}^{-1}$, and $\kappa_{b,c}$, the internal loss and coupling rates, respectively. $\kappa_{b,c}/2\pi$, is shown in Fig. 2c as a function of the detuning from the Purcell, $\omega_{b}-\omega_{pb}$. A large tuning range is observed, ranging from $\sim 10$ kHz up to $\sim 3$ MHz. The tuning curve is identical for all values of buffer resonator frequency (see Fig. 2c). The buffer is over-coupled in a large parameter range (corresponding to $\kappa_{b,c}\gg\kappa_{b,i}$), which is desirable for SMPD operation, as buffer internal losses further reduce $\eta_{\mathrm{4WM}}$ [1].
Qubit readout is performed at the waste resonator frequency when the qubit is in the excited state, $\omega_{W}-\chi_{W}$, to minimize the chances to excite it out of the ground state and cause unwanted dark counts, with a readout duration $T_{\mathrm{RO}}=0.8$ µs. Readout histograms are shown in Fig. 3a, yielding the readout fidelity $\mathcal{F}_{\mathrm{RO}}=0.87$ and equilibrium excited state population $p_{\mathrm{th},q}=8.5\cdot 10^{-4}$. The qubit lifetime is $T_{1}=70$ µs (see Fig. 3b), and the coherence time measured under a Ramsey sequence is $T_{2}^{*}=20$ µs. The dispersive shifts $\chi_{b}$ and $\chi_{w}$ are reported in Table 2. Active reset is performed as explained in Section IIC [1].
Figure 3: a) Qubit readout histograms after 95650 sequences, in absence of any preparation pulse (blue full circles), and after a $\pi$ pulse (orange full circles). Solid lines are fits by a sum of two Gaussians centered on the two quadrature $I$ values corresponding to the ground and excited states, yielding $p_{q,th}=8.5\cdot 10^{-4}$, and $\mathcal{F}_{\mathrm{RO}}=0.87$. The red line shows the threshold. b) Qubit excited state probability following a $\pi$ pulse, yielding $T_{1}=70$ µs. c) Qubit Ramsey fringes yielding $T_{2}^{*}=20$ µs. Insets : corresponding pulse sequences. Purple pulses are sent through the drive line, green through the readout (RO) line.

### v2

# Louis Pallegoix1, Jaime Travesedo1, Alexandre S. May1,2, Léo Balembois1, Denis Vion1, Patrice Bertet*1 and Emmanuel Flurin Université Paris-Saclay, CEA, CNRS, SPEC, 91191 Gif-sur-Yvette Cedex, France 2Alice&\&Bob, 53 boulevard du Général Martial Valin, 75015 Paris

@@ Abstract
We report on the characteristics of a microwave photon counter device based on a superconducting transmon qubit. Its design is similar to [4], with an additional bandwidth tuning circuit that permits tuning the detection bandwidth over an order of magnitude to optimize the device efficiency and noise. Owing to this new feature and to improvements in device fabrication, a power sensitivity of $3\cdot 10^{-23}\ \mathrm{W}/\sqrt{\mathrm{Hz}}$ is reached. We confirm the high performance of the device by measuring single spin microwave fluorescence.
@@ I Introduction
Single-photon detection has long been a fundamental tool in quantum technologies, with well-established applications in the optical domain, ranging from fluorescence microscopy [32, 22, 9, 14] to quantum communication [18] and boson sampling [5] ; however, extending single-photon detection into the microwave regime presents significant challenges. The energy of microwave photons—of the order of µeV—is several orders of magnitude lower than that of optical photons, requiring detectors to operate at cryogenic temperatures to suppress thermal noise and reduce the background thermal photon population.
Despite these challenges, single-microwave-photon detectors (SMPDs) have emerged as critical tools for advancing quantum sensing and information processing. Their ability to detect individual microwave photons has opened possibilities for a wide range of applications, including the detection of weak incoherent microwave emitters, such as electron spins in solids [1, 41, 10], and even searches for hypothetical dark-matter particles [26, 16]. SMPDs also hold potential for primary thermometry at low temperatures [38] and for use in quantum illumination protocols [3]. Moreover, these detectors can play a key role in quantum information processing [33, 13, 6], enabling heralded entanglement between superconducting qubits [29], improving qubit readout schemes [31], and facilitating the generation of robust quantum states [7].
Various SMPD designs have been proposed and implemented, with most relying on superconducting qubits [28, 1, 29, 34, 20, 36, 25, 37, 19, 35, 15, 24, 21, 8, 23], quantum dots [42, 17] or bolometric detectors [27]. These devices have demonstrated promising results, achieving sensitivities ranging from $10^{-21}$ to $10^{-22}\,\mathrm{W/\sqrt{Hz}}$ at microwave frequencies ; however, there remain limitations in terms of bandwidth, detection efficiency, and dark count rates that must be addressed to further enhance their performance.
[…]
@@ II.1 Device design
The SMPD device (see Figs. 1a and b) consists of a transmon qubit (mode $q$, frequency $\omega_{q}$), capacitively coupled to two harmonic oscillator modes, called the buffer (mode $b$, frequency $\omega_{b}$), and the waste (mode $w$, frequency $\omega_{w}$). The buffer is connected to the input line of the device through another resonator at frequency $\omega_{pb}$. The latter serves as Purcell filter to decouple the qubit from the input line, while also setting the effective buffer energy coupling rate $\kappa_{\mathrm{b,c}}$. The total energy-loss rate of the buffer resonator is $\kappa_{b}=\kappa_{\mathrm{b,c}}+\kappa_{\mathrm{b,i}}$, where $\kappa_{\mathrm{b,i}}$ represents the internal losses of the buffer. The waste is connected to an output line through another Purcell filter, at frequency $\omega_{pw}$, with a rate $\kappa_{w}$. Other relevant parameters are the qubit energy-relaxation time $T_{1}$ and equilibrium excited-state population, $p_{\mathrm{th},q}$, which varies with the qubit temperature $T$, in the $T\ll\hbar\omega/k_{B}$ limit, as $p_{\mathrm{th},q}=\frac{1}{e^{\frac{\hbar\omega_{q}}{k_{B}T}}-1}$. In this formula, $T$ is not necessarily the cryostat base temperature, as various heating effects (due to infrared radiation in particular) may lead to a higher effective qubit temperature.
For optimal operation as an energy detector, both the frequency and bandwidth of the device need to be matched to the source of interest, be it a spin [4, 41] or a dark-matter candidate [12]. Frequency tuning is achieved by inserting a superconducting quantum interference device (SQUID) into the buffer mode, resulting in a buffer frequency $\omega_{b}(\phi_{b})$ that depends on the flux threading the buffer SQUID loop, $\phi_{b}$. An asymmetric SQUID is purposely chosen to minimize sensitivity to flux noise. Bandwidth tuning is achieved by inserting a second SQUID into the Purcell filter mode, resulting in a buffer coupling rate $\kappa_{\mathrm{b,c}}(\phi_{\mathrm{pb}})$ that depends on the flux threading the Purcell SQUID loop, $\phi_{\mathrm{pb}}$. This SQUID is symmetric to maximize the bandwidth-tuning range.
@@ II.2 Four-wave mixing
The working principle of photon detection relies on a 4WM process that maps the presence or absence of a photon onto the state of the transmon qubit. The nonlinear element is provided by the qubit Josephson junction. The 4WM is activated by a microwave signal (called the pump, frequency $\omega_{p}$) of amplitude $\xi_{0}$, applied to the qubit via a pump line connected capacitively to the qubit (see Fig. 1). When the pump frequency approximately satisfies the relation $\omega_{b}+\omega_{p}=\omega_{q}+\omega_{w}-\chi_{w}$, the dominant process is described by the Hamiltonian
@@
where $\chi_{b}$ (resp. $\chi_{w}$) is the qubit dispersive coupling to the buffer mode (resp. the waste mode) [28]. This Hamiltonian reversibly converts a buffer and pump photon into a qubit excitation and waste photon. Dissipation of the waste photon into the output line ensures that the process is irreversible ; nonetheless, the qubit remains excited long enough to be readout subsequently, and it can thus be used as a flag that frequency conversion occurred.
An approximate description is obtained by eliminating the qubit degree of freedom, yielding an effective Hamiltonian $H_{\mathrm{eff}}=g_{\mathrm{4WM}}bw^{\dagger}+h.c.$ with $g_{\mathrm{4WM}}=\sqrt{\chi_{b}\chi_{w}}\xi_{0}$. This describes the parametric coupling between the buffer and waste modes, similar to the Josephson parametric converter for instance. Assuming that $\kappa_{\mathrm{b,i}}\ll\kappa_{\mathrm{b,c}}$, the efficiency of frequency conversion under this process is $\eta_{\mathrm{4WM}}=4C/(1+C)^{2}$, where $C=\frac{4\chi_{b}\chi_{w}|\xi_{0}|^{2}}{\kappa_{w}\kappa_{b}}$ is the system cooperativity. Unit efficiency is reached when $C=1$. Physically, this occurs when the frequency-conversion rate $4|g_{\mathrm{4WM}}|^{2}/\kappa_{w}$ is equal to the input resonator coupling rate, $\kappa_{b}$. This condition can always be achieved by adjusting the pump frequency and amplitude (see Appendix A for more details).
[…]
@@ II.3 Cyclic operation
[…]
In the third step, the qubit is reset into its ground state by applying a $\pi$ pulse conditioned to the outcome of the previous readout being the excited state. This efficiently resets the qubit into its ground state, with a probability that can be even lower than the thermal-equilibrium qubit excited-state probability [4]. Note that the duration of the reset step $T_{\mathrm{reset}}$ varies from cycle to cycle due to its conditional nature. The detector duty cycle $\eta_{\mathrm{cycle}}$ is defined as the average value of $T_{d}/(T_{d}+T_{\mathrm{RO}}+T_{\mathrm{reset}})$.
@@ III Figures of merit
In this section, we describe in some detail the processes that limit the performance of the device as an energy detector. This performance is captured by the power sensitivity $S$, defined as the power that can be detected with unit signal-to-noise ratio in $1$ s integration time [4]:
@@
In this formula, $\eta_{\mathrm{SMPD}}$ is the overall detection efficiency of the SMPD and $\alpha$ is the average number of counts per second when no signal is sent to the device. The latter can be separated in two contributions $\alpha=\alpha_{\mathrm{err}}+\alpha_{\mathrm{th}}$, $\alpha_{\mathrm{err}}$ being the false-positive events and $\alpha_{\mathrm{th}}$ being the counts due the detection of thermal photons in the input line. In what follows, we address each of these three quantities, with the goal of maximizing $\eta_{\mathrm{SMPD}}$ while minimizing $\alpha_{\mathrm{err}}$ and $\alpha_{\mathrm{th}}$.
@@ III.1 Detection efficiency
[…]
where the detector frequency response function is approximated by a Lorentzian of width $\kappa_{d}$ (the detector linewidth) such that $\eta_{\omega}=\frac{1}{1+\left[\frac{2(\omega-\omega_{b})}{\kappa_{d}}\right]^{2}}$ . The qubit efficiency $\eta_{q}=\frac{T_{1}}{T_{d}}(1-\mathrm{e}^{-\frac{T_{d}}{T_{1}}})$ can be approximated by $1-\frac{T_{d}}{2T_{1}}$ in the limit $T_{d}\ll T_{1}$ [4]. To maximize $\eta_{\mathrm{SMPD}}$, $T_{d}$ should be optimized to maximize the duty cycle while keeping the cycle time well below $T_{1}$ to avoid relaxation of the qubit before readout.
@@ III.2 Dark count rate
The dark count rate $\alpha_{\mathrm{err}}$ is the number of counts detected per unit of time, even though no photon was present in the input line. It has two contributions, $\alpha_{\mathrm{err}}=\alpha_{q}+\alpha_{p}$. First, the qubit can be found in the excited state despite the active reset procedure. This is due to the relaxation toward the thermal-equilibrium value $p_{q,th}$, which takes place during the detection window with a characteristic time $T_{1}$ [4], giving rise to $\alpha_{q}=p_{th,q}(T_{d}/T_{1})(1/(T_{d}+T_{\mathrm{RO}}+T_{\mathrm{reset}}))$ (assuming that the active reset works perfectly, otherwise an offset must be added to that expression). The other contribution $\alpha_{p}$ is due to heating brought by the pump. Note that the readout errors causing the ground state to be erroneously detected as the excited state are negligible because we purposely choose a detection threshold that minimizes this effect.
@@ III.3 Thermal counts
Thermalizing the microwave field at millikelvin temperatures is well known to be a difficult task because of vanishing thermal conductivity and the large attenuation needed in the input lines. This results in an effective field temperature that is often significantly larger than the cryostat base temperature, and in our case, in the presence of a finite flux of thermal photons at the SMPD input, causing a thermal count rate of thermal origin
$\alpha_{\mathrm{th}}=\frac{1}{2\pi}\int_{-\infty}^{\infty}\bar{n}_{\mathrm{th,b}}\eta_{\mathrm{SMPD}}(\omega)d\omega,$ (4)
where $\bar{n}_{\mathrm{th,b}}=\frac{1}{e^{\frac{\hbar\omega_{b}}{k_{B}T}}-1}$ is the number of photons per mode assuming that the field is at temperature $T$. Because $\kappa_{d}\ll\omega_{b}$, we get
@@
Due to the tunable bandwidth of our device, we are able here to test the predicted proportionality between noise and detector bandwidth, as will be seen below.
@@ IV.1 Device characterization
The SMPD resonators are measured with a vector network analyser (VNA). Each resonance is fitted to determine the resonator frequency and linewidth. The results of these fits are reported in Tab. 1.
@@
Table 1: Measured resonator frequencies and linewidths. The Purcell filter of the buffer and the buffer itself are tunable over the indicated interval. The buffer linewidth depends on the the Purcell frequency, as shown in fig. 2.
@@
Figure 2: Tunability curves. a) Solid circles are the measured frequency of the buffer resonator, $\omega_{b}$, as a function of the flux bias applied to its SQUID, $\Phi_{b}$, for several values of $\Phi_{pb}$ (blue is $0.45\Phi_{0}$, orange is $0.24\Phi_{0}$, green is $0.03\Phi_{0}$). The solid lines are sinusoidal fits, yielding the asymmetry between the buffer SQUID junctions. b) Full blue circles are the measured frequency of the Purcell filter, $\omega_{pb}$, as a function of the flux bias sent to its SQUID $\Phi_{pb}$. The solid line is a fit assuming a symmetric SQUID. c) Full circles show the measured coupling rate of the buffer resonator to the $50\Omega$ line, as a function of its detuning from the Purcell filter, $\omega_{b}-\omega_{pb}$, for 33 different values of $\Phi_{b}$ : blue for $\Phi_{b}=0$, orange for $\Phi_{b}=0.25\Phi_{0}$, green for $\Phi_{b}=0.5\Phi_{0}$. The dashed line shows the internal losses of the buffer, $\kappa_{b,i}$. The solid lines are lorentzian fits to data of the same color. We can see that $\kappa_{b}/2\pi$ rise above 3 MHz at zero detuning and fall below $\kappa_{b,i}$ when the detuning exceeds 100 MHz in absolute value. d) Full circles are the measured SMPD efficiency, $\eta_{\mathrm{SMPD}}$, as a function of signal frequency $\omega$, for $\kappa_{b}/2\pi=170$ kHz (blue) and $280$ kHz (orange). Dashed lines are Lorentzian fits yielding the SMPD bandwidth $\kappa_{d}/2\pi$, equal to $314$ kHz and $531$ kHz, respectively.
@@
Figure 2a shows $\omega_{b}(\phi_{b})$ for several values of $\phi_{pb}$. The buffer tuning range is narrow (60 MHz) and the curve is well fitted by a sinusoid [11], from which we deduce a ratio of 15 between the Josephson-junction critical currents of the asymmetric SQUID. Little crosstalk between the $\phi_{b}$ and $\phi_{pb}$ is observed, as evidenced by the weak dependence of $\omega_{b}/2\pi$ on $\phi_{pb}$.
@@
From the fits of the buffer resonance curves, we extract $\kappa_{b,i}=2.2\cdot 10^{5}\mathrm{s}^{-1}$, and $\kappa_{b,c}$, the internal loss and coupling rates, respectively. Fig. 2c shows $\kappa_{b,c}/2\pi$ as a function of the detuning from the Purcell, $\omega_{b}-\omega_{pb}$. A large tuning range is observed, ranging from $\sim 10$ kHz up to $\sim 3$ MHz. The tuning curve is identical for all values of buffer-resonator frequency (see Fig. 2c). The buffer is overcoupled over a large parameter range (corresponding to $\kappa_{b,c}\gg\kappa_{b,i}$), which is desirable for SMPD operation, as buffer internal losses further reduce $\eta_{\mathrm{4WM}}$ [4].
[…]